OpenStudy (anonymous):

a number is six times another number. when the numbers are added their sum is 49. what are the numbers

OpenStudy (lgbasallote):

a number is six times another number let x = a number y = another number therefore x = 6y their sum is 49 x + y = 49 if i substitute x into 6y 6y + y = 49 7y = 49 y = 1 can you find x now?

OpenStudy (anonymous):

I know that 6 x 7 + 42 = 49 but I don't know how to show my work on it.

OpenStudy (lgbasallote):

the work is what i did...you just continue

OpenStudy (lgbasallote):

btw... i wrote the final answer wrong\ should be y = 7 hehe

OpenStudy (lgbasallote):

maybe that's why you were confused

OpenStudy (anonymous):

yeah maybe. Thank you for your help. :)

OpenStudy (lgbasallote):

welcome welcome :d

OpenStudy (lgbasallote):

:D *